Find the roots. 2x^2+13x+15=0
Found 2 solutions by nerdybill, richwmiller:
Answer by nerdybill(7384)

2x^2+13x+15=0
rewrite the middle term as:
2x^2+10x+3x+15=0
group terms:
(2x^2+10x)+(3x+15)=0
factor each group:
2x(x+5)+3(x+5)=0
(x+5)(2x+3) = 0
x = {-5, -3/2}

Answer by richwmiller(17219)

find factors of 30 that add up to 13
10 and 3
2x^2+10x+3x+15
2x*(x+5)+3(x+5)
(2x+3)*(x+5)
2x+3=0
2x=-3
x=-3/2
x+5=0
x=-5